find x+1/x if x-1/x= square root of 5
Found 2 solutions by ptaylor, mathstutor494:
Answer by ptaylor(2198)

x-1/x=sqrt5
let's add 2/x to each side and we get
x-1/x +2/x=sqrt5+2/x simplify
x+1/x=sqrt5+2/x=(xsqrt5+2)/x
Hope this helps---ptaylor

Answer by mathstutor494(120)

x-1/x = square root of 5
Squaring both sides
>x^2+(1/x^2)-2=5
>x^2+(1/x^2)=5+2
>x^2+(1/x^2)=7
Adding 2 to both sides
>x^2+(1/x^2)+2=7+2
>(x+1/x)^2= 3^2
Square root of above eqn...
Ans x+1/x=3